At a store that has 1500 CD's and 980 DVD's in stock, 10 of the CD's and 4 of the DVD's are known to have a scratch on them. If a customer picks a disk at random, what is the probability that it will be

(a) a DVD or scratched?

(b) a CD or not scratched?

(c) a DVD or a CD?

Explain work.
Answer #1

A) P(DVD or scratch) = 990/2480 =0.399

B) P( CD or not scratch ) = 2476/2480 = 0.998

C) P( CD or DVD) = 1

Solution file is attached go through it

Soll At a store that hes 1500 CDs and 980 Dvds in stock; lo of CDs and 4 of DVDs are known to love scratch on them. If Cu

Thanks